OpenStudy (moongazer):

solve for the real numbers X and Y 1.) 5i=x-yi 2.) x-3yi=8

OpenStudy (moongazer):

more to come ill back later for clarifications

OpenStudy (anonymous):

1) => y = -5 (if it should be real) but than 2) is not solve able in 2) y must be something like 0

OpenStudy (anonymous):

Last time I checked zero was still a real number.

OpenStudy (anonymous):

but the answers does not match 1) x := 0 , y := -5 2) x := 8 , y := 0 the x and the y are not the same !

OpenStudy (anonymous):

We'll have to wait for moongazer to return to clarify if this is a system of equations or not.

OpenStudy (anonymous):

Get everything over to the same side of the equation: 0 = x + 5 - y)i = 0 + 0i. An important property of complex numbers is that two of them are equal if and only if their real components arr equal, and their complex components are equal. With that in mind, we have the following: (i) x = 0, and (ii) 5 - y = 0, implying y = 5.

OpenStudy (moongazer):

I'm back! so, how do we start?

OpenStudy (anonymous):

Start by getting everything over to the same side of the equation.

OpenStudy (moongazer):

=x-yi=5i like that?

OpenStudy (anonymous):

Have 0 be on one side. Everything else on the other.

OpenStudy (moongazer):

0=x-yi-5i? why should there be a zero if i am correct?

OpenStudy (anonymous):

Now group all terms together that have an i in them.

OpenStudy (moongazer):

why should there be a zero? 0=x-(y-5)i ?

OpenStudy (anonymous):

An important thing to remember about complex numbers is that two of them are defined to be equal if and only if their real parts are equal, and their imaginary parts are equal.

OpenStudy (anonymous):

The number 0 can be written as 0 + 0i.

OpenStudy (anonymous):

A post appeared about twenty minutes that outlines the procedure that determines x = 0 and y = 5.

OpenStudy (anonymous):

...twenty minutes ago...

OpenStudy (moongazer):

sorry for the late reply

OpenStudy (moongazer):

what happened to i?

OpenStudy (anonymous):

Did you group the terms that have an i in them?

OpenStudy (anonymous):

You might have x + (5 - y)i = 0 + 0i.

OpenStudy (anonymous):

The real part of the left side of the equation is x, and its imaginary part is 5 - y,

OpenStudy (anonymous):

The right-hand side of the equation has 0 as both its real and imaginary parts.

OpenStudy (anonymous):

All that's left to do is set the corresponding parts equal.

OpenStudy (moongazer):

what is the meaning of if and only if?

OpenStudy (anonymous):

You'll have x = 0, and 5 - y = 0, implying y = 5.

OpenStudy (anonymous):

If and only if mrans that the implication goes both ways. "A if and only if B" means that A implies B, and B implies A.

OpenStudy (moongazer):

so the steps are ; put the equation on one side like =x-(y-5)i, and equate it to 0+0i ? so it will become 0+0i = x-(y-5)i then divide all sides by i so it will become 0= x-y-5 is this correct?

OpenStudy (anonymous):

You got the gist of it. Concerning the real and imaginary parts of a complex number: they are real numbers, denoted bt Re(z) anf Im(z), where z is a complex number.

OpenStudy (anonymous):

... by Re(z) and Im(z)...

OpenStudy (moongazer):

please explain it in an easy manner i don't know what " denoted bt Re(z) anf Im(z), where z is a complex number" means

OpenStudy (moongazer):

so, if my explanation is correct how did you get x and y? sorry if i can't understand it easily.

OpenStudy (moongazer):

could you make a conclusion that can easily be understand?

OpenStudy (anonymous):

How do you tell if two complex numbers are equal?

OpenStudy (moongazer):

if the imaginary and the real numbers have the same value

OpenStudy (anonymous):

Right. So if I say that 2 + 3i = 2 - 5i, am I right?

OpenStudy (moongazer):

no

OpenStudy (anonymous):

You get the golden goode award :^)

OpenStudy (anonymous):

...goose award.

OpenStudy (moongazer):

?

OpenStudy (anonymous):

Just goofing. You are right.

OpenStudy (moongazer):

then what now? ^_^

OpenStudy (anonymous):

Do you see why x = 0?

OpenStudy (moongazer):

no i am stuck with x-y-5

OpenStudy (moongazer):

=0

OpenStudy (anonymous):

You should have x + (5 - y)i = 0 + 0i.

OpenStudy (moongazer):

then?

OpenStudy (anonymous):

The i is still there. It doesn't vanish.

OpenStudy (moongazer):

so, how do you equate it?

OpenStudy (anonymous):

If the complex number on the left is the same as 0 + 0i, then x = 0, and 5 - y = 0.

OpenStudy (anonymous):

Solve that last equation for y.